OpenStudy (anonymous):

Solve the problem. Devon purchased tickets to an air show for 4 adults and 2 children. The total cost was . The cost of a child's ticket was less than the cost of an adult's ticket. Find the price of an adult's ticket and a child's ticket.

OpenStudy (anonymous):

x=$4 ans. for the price of the children tickets. 4+8=$12 ans. for the price of the adult tickets. Proof: 7*12+2*4=92 84+8=92 92=92

OpenStudy (anonymous):

these are the multiple choice answers: adult's ticket: $13; child's ticket: $10 adult's ticket: $12; child's ticket: $9 adult's ticket: $14; child's ticket: $11 adult's ticket: $11; child's ticket: $8

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Thank you for helping ^^

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

wait what was the total cost for all the tickets

OpenStudy (anonymous):

$66

OpenStudy (anonymous):

ok so this might be the long way but this is how i got your answer

OpenStudy (anonymous):

so your answer chocies consist of this 13,10 and 12,9 and 14,11 and 11,8

OpenStudy (anonymous):

ok o :

OpenStudy (anonymous):

it says that adult has to cost more than the child ticket so therefore all of the big numbers are adult tickets and all of the small are kids. what i did is i saw that there where 4 adults and 2 kids so i got on my calculator :D and i did this.\[12\times4=48 + 9\times2=18\]

OpenStudy (anonymous):

if u add 48 plus 18 u get the answer of 66

OpenStudy (anonymous):

ohhh o.o lol....

OpenStudy (anonymous):

um... did you use the answer choices to make that equation? o.o

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

your answer is adult's ticket: $12; child's ticket: $9

OpenStudy (anonymous):

I am talking about 12 * 4 = 48 + 9 *2=18

OpenStudy (anonymous):

Sorry T:

OpenStudy (anonymous):

o ya i got that from the answer choices

OpenStudy (anonymous):

you can do that? T.T;

OpenStudy (anonymous):

Thank you for your help once again ^^

OpenStudy (anonymous):

lol your confusing the crap out of me lol.. ok what do u mean i can do that

OpenStudy (anonymous):

I'm sorry, nevermind T_T

OpenStudy (anonymous):

lol no im here to help math is my second strongest subject so throw it at me

OpenStudy (anonymous):

Then, I have to throw a lot of questions at you... LOL D:

OpenStudy (anonymous):

o god lol but ok.. what grade are u in??

OpenStudy (anonymous):

I'm a Freshmen in college.. T:

OpenStudy (anonymous):

not good at math so I'm starting with basic level.

OpenStudy (anonymous):

lol my sister is a freshman in college im in 11th grade

OpenStudy (anonymous):

oh ^^; i feel embarrassed now..

OpenStudy (anonymous):

haha dont be :D i like helping out people :)

OpenStudy (anonymous):

and im like u im taking basic Spanish so im in class with a lot of freshman and 10th graders

OpenStudy (anonymous):

oh o :

OpenStudy (anonymous):

ya so ask me your questions

OpenStudy (anonymous):

Without solving, choose the correct solution by deciding which choice satisfies the given conditions. At the local convenience store, 3 bags of chips and 7 containers of dip cost $31. However, 6 bags of chips and 3 containers of dip cost $18. What is the cost of one bag of chips and one container of dip? chips = $3; dip = $5 chips = $1; dip = $4 chips = $1.50; dip = $3.50 chips = $2; dip = $3

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

ok lets do this

OpenStudy (anonymous):

lol :(

OpenStudy (anonymous):

wdf hold on

OpenStudy (anonymous):

lol sorry im laughing a lot i cant think right now

OpenStudy (anonymous):

its ok...

OpenStudy (anonymous):

y are u so nice???

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

lol sorry i was just wondering

OpenStudy (anonymous):

chips = $1; dip = $4

OpenStudy (anonymous):

thats your answer u want me to explain or no??

OpenStudy (anonymous):

explain please o:

OpenStudy (anonymous):

ok so we know that the answer choice has to be correct for both problems instead of just one another

OpenStudy (anonymous):

so what i did was i looked at you answers and i did this it says At the local convenience store, 3 bags of chips and 7 containers of dip cost $31. However, 6 bags of chips and 3 containers of dip cost $18. i looked at the answer choices chips = $3; dip = $5 chips = $1; dip = $4 chips = $1.50; dip = $3.50 chips = $2; dip = $3 and i looked at the first one and it says dip= $5 and i looked at your first problem 3 bags of chips and 7 containers of dip cost $31. and i multiplied 5times7 and got 35 see now i know that the first answer is wrong because it already goes over what its suppose to equal witch is 31$

OpenStudy (anonymous):

then i did the second answer witch is chips = $1; dip = $4 so i got the dip amount witch is 4 and i multiplied that by 7 witch is 28 then is looked at how much the chips cost witch is 1 dollar per bag and it says that it has 3 bags so i added 28 plus 3 and u get 31 so that means answer 2 is correct for the 1st problem. for the second problem 6 bags of chips and 3 containers of dip cost $18. i did the same thing i did with the first problem i multiplied 4times3 equals 12 plus 6=18 and therefore your answer for this problem is chips = $1; dip = $4

OpenStudy (anonymous):

sorry if its long

OpenStudy (anonymous):

No, it's not long. It's fine. I understand now :)

OpenStudy (anonymous):

:D yay any more questions

OpenStudy (anonymous):

I could keep asking you? ^^; hope it doesn't bother you.

OpenStudy (anonymous):

lol ya u can keep asking

OpenStudy (anonymous):

okay ^^.

OpenStudy (anonymous):

Solve the system of equations by the addition method. 3x+7y=-9 -5x-3y=-11

OpenStudy (anonymous):

Answer Choices: (4, 3) (-4, -3) (4, -3) (-4, 3)

OpenStudy (anonymous):

O.o i always hated doing this.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

lol im still gonna help u tho :D

OpenStudy (anonymous):

Thank you :3

OpenStudy (anonymous):

man i feel dumb right now

OpenStudy (anonymous):

Take your time and if you don't know its okay o:

OpenStudy (anonymous):

It's okay I got 24/30 on the quiz. O: thank you for your help hehe.

OpenStudy (anonymous):

:) yay u got an 80 :DDDDDDDDDDDDDD

OpenStudy (anonymous):

:( lol not good enough!